# Traffic Engineering in Large-scale Networks with Generalizable Graph Neural Networks

### v1

@@ IV-B IPM Preliminaries
The primary bottleneck of IPM is the complexity of solving large linear systems of equations (the Newton’s step in Line 5) at each iteration. The fastest linear solver runs in almost cubic time, significantly limiting scalability of IPM beyond modest problem sizes. Meanwhile, due to its use of the Newton’s step, IPM can converge to optimality in almost a constant number of iterations in most practical settings, typically less than $5$-$30$ iterations regardless of problem size [15]. This makes IPM a particularly good candidate to be mimicked and sped-up by ML—speeding-up the Newton’s step will yield good asymptotic improvement, while the small number of iterations is ideal for algorithmic alignment, as we discuss in Sec. IV-D.
@@ IV-C Representing TE as a Graph
To speed-up TE, our first step is to represent the TE problem as a neural model that can be trained. As we motivated in Sec. IV-A, a simple GNN defined over the network topology graph itself cannot effectively embed information such as decision variables (path traffic split ratios), constraints and the objective. To this end, we develop a new GNN architecture, where the graph is not defined on the network but on all the entities (variables, constraints and objective) and their correlations presented in the LP formulation in Program (6).
@@
The key principle behind the above construction, as also hinted by existing work [39, 6, 32, 20], is to fully represent all information within the LP formulation in the same model. Correlations among variables, constraints and the objective are fully captured by the edges, while all problem data are encoded as weights on the edges. The model can further easily generalize to adding, removing or permuting of variables and constraints. Next, we further talk about how the GNN model can be aligned to the IPM algorithm, and how it can also be trained under strong supervision by IPM.
Figure 2: Double-looped GNN architecture: $K$ outer loops simulating IPM iterations, and $J$ inner layers to learn an IPM step.
@@ IV-D Algorithmic Alignment and Strongly Supervised Training
Motivated by the algorithmic alignment literature in ML [6, 38], we are inspired to better align our GNN model with the IPM algorithm to achieve strong performance and generalizability. Since an IPM iteration basically takes a Newton’s step from the previous iteration’s solution, starting from an initial feasible solution, we want the GNN model to learn the same behavior. To this end, we define a double-looped GNN architecture, which consists of $K$ outer GNN layers, and $J$ inner GNN layers in each outer layer, as in Fig. 2. Each outer layer corresponds to an IPM iteration, and seeks to mimic one Newton’s update on the previous layer’s solution (represented by vertex attributes after previous layer). The number of outer layers, $K$, should correspond to the number of iterations that IPM takes to converge, which is usually a constant that can be easily estimated [32].
Because each outer layer corresponds to the same operation (Newton’s step) that is applied on a different input (previous iteration’s solution), the message passing function ${\sf msg}^{k}$ and the aggregation functions ${\sf aggr}^{k}$ in all outer GNN layers should share the same neural network parameters. Furthermore, since the IPM generates a feasible solution in every iteration, each GNN outer layer’s output can be supervised by a loss function defined between the GNN’s output solution (by applying the readout function $r$) and the IPM’s updated solution. This strong supervision of each GNN layer’s output can greatly improve the model’s ability to learn not just the optimal solutions in the training dataset, but also the trajectories of convergence to optimality by IPM. This is crucial for the neural model to learn the generalizability of the optimal algorithm—to converge to optimality given any problem instance.
One concern in the above architecture is whether one (outer) GNN layer can well approximate the Newton’s step, which is a fairly complex operation that warrants a nearly cubic complexity. To this end, we further expand each outer layer to consist of $J$ inner GNN layers. Intuitively, these layers are to learn the intermediate steps in solving the Newton’s system, plus applying the system’s solution to the variables’ update. Since these steps do not correspond to feasible solutions to the original problem, the inner GNN layers are unsupervised. In practice, we find that $J=2$ inner layers are already sufficient for even the largest networks that we evaluate.
@@ IV-E Model Design and Training Procedures
With the above discussions, we now describe the GNN model and training procedures for solving large-scale TE problems with generalizability as depicted in Fig. 2. The GNN model consists of $J$ sets of message passing and aggregation functions $\bm{{\sf msg}}_{j}$ and $\bm{{\sf aggr}}_{j}$ corresponding to the $J$ inner layers, and one readout function $r$. Each $\bm{{\sf msg}}_{j}$ contains six distinct functions: ${{\sf msg}}_{j}^{\sf p\text{-}dl}$, ${{\sf msg}}_{j}^{\sf dl\text{-}p}$, ${{\sf msg}}_{j}^{\sf p\text{-}o}$, ${{\sf msg}}_{j}^{\sf o\text{-}p}$, ${{\sf msg}}_{j}^{\sf dl\text{-}o}$, and ${{\sf msg}}_{j}^{\sf o\text{-}dl}$. Each $\bm{{\sf aggr}}_{j}$ contains three functions: ${\sf aggr}_{j}^{\sf p}$, ${\sf aggr}_{j}^{\sf dl}$, and ${\sf aggr}_{j}^{\sf o}$. For $\alpha,\beta\in\{{\sf p,dl,o}\}$, the functions ${{\sf msg}}_{j}^{\sf\alpha\text{-}\beta}$ and ${\sf aggr}_{j}^{\sf\beta}$ are used to update the vertex attributes of $V^{\beta}$ based on the attributes of $V^{\alpha}$, where ${\sf p}$ denotes path vertices, ${\sf dl}$ denotes demand and link vertices, and ${\sf o}$ denotes the objective vertex. We cluster demand and link vertices together since all constraints can be normalized to have the same right-hand-side, i.e., let $\mathbf{b}=1$ in the LP formulation. All functions are shared across all outer layers, and the model can have a variable number of outer layers depending on the input instance.
@@
$h_{\text{dl}}^{k,j}=$ ${\sf aggr}_{j}^{\text{dl}}\Big{(}h_{\text{dl}}^{k,j-1},{\sf msg}_{j}^{\text{p-dl}}(\{h_{\text{p}}^{k,j-1}|e_{p,i}\in E^{\text{pd}},e_{p,l}\in E^{\text{pl}}\}),$
${\sf msg}_{j}^{\text{o-dl}}(\{h_{\text{o}}^{k,j-1}|e_{i,o}\in E^{\text{do}},e_{l,o}\in E^{\text{lo}}\})\Big{)};$
$h_{\text{o}}^{k,j}=$ ${\sf aggr}_{j}^{\text{o}}\Big{(}h_{\text{o}}^{k,j-1},{\sf msg}_{j}^{\text{dl-o}}(\{h_{\text{dl}}^{k,j}|e_{i,o}\in E^{\text{do}},e_{l,o}\in E^{\text{lo}}\}),$
${\sf msg}_{j}^{\text{p-o}}(\{h_{\text{p}}^{k,j-1}|e_{p,o}\in E^{\text{po}}\Big{)};$
$h_{\text{p}}^{k,j}=$ ${\sf aggr}_{j}^{\text{p}}\Big{(}h_{\text{p}}^{k,j-1},{\sf msg}_{j}^{\text{dl-p}}(\{h_{\text{dl}}^{k,j}|e_{p,i}\in E^{\text{pd}},e_{p,l}\in E^{\text{pl}}\}),$
${\sf msg}_{j}^{\text{o-p}}(\{h_{\text{o}}^{k,j}|e_{p,o}\in E^{\text{po}}\})\})\Big{)}.$
@@
where ReLU is a rectified linear unit, and $D^{(\tau)}$ and $L^{(\tau)}$ are the corresponding demand set and link set in problem instance $\tau$. The constraint loss is defined as the constraint violation of solution generated in every GNN outer loop, normalized by the right-hand-side constant of each constraint. It penalizes the output solution of each outer loop to ensure feasibility of the solution update trajectory, in line with the original IPM.
@@
The objective loss penalizes when the generated solutions deviate too much from the IPM’s updates in each iteration.
@@
$\mathcal{L}=\rho_{1}\mathcal{L}_{\sf p}+\rho_{2}\mathcal{L}_{\sf dl}+\rho_{3}\mathcal{L}_{\sf o}.$ (3)
@@ V-A Datasets
Erdös-Rényi (ER) Dataset: We generated random topologies based on the Erdös-Rényi model [10]. For training set, we generated networks with $\{20,30,\dots,100\}$ nodes, each with a connection probability of $\{0.3,0.4,\dots,0.8\}$. This resulted in $9\times 6=54$ topology configurations, and we generated $200$ instances for each configuration. For each instance, $10$ random SD pairs were selected. The link capacities and SD pair demands were both randomly generated in $[1000,5000]$; the same default range applies to all other datasets below. For test set, we generated significantly larger networks with $\{200,500,1000,2000\}$ nodes respectively, with connection probability in $\{0.7,0.8,0.9\}$. Each of the $12$ test configuration had $300$ instances with the same number of SD pairs and range for demands and capacities; we denote these $12$ test sets as $\text{ER}^{\sf test}_{n,q}$, where $n$ is the node number, $q$ is the connection probability. The largest topology in training set had almost $8\times 10^{3}$ links, and the largest topology in test set had $3.6\times 10^{6}$ links, two to three orders of magnitude larger than training.
@@
Waxman (WA) Dataset: We generated random networks based on Waxman’s model [36]. For training set, we generated networks with $\{200,300,\dots,800\}$ nodes with default connection parameters $\alpha=0.1$ and $\beta=\{0.2,0.18,\dots,0.08\}$ [^1] Here the $\alpha,\beta$ are Waxman parameters [36], not those used in Sec. IV-E.. Each configuration had $800$ instances, and each instance had $20$ random SD pairs. For test set, we generated large networks that had $\{2000,3000,4000,5000\}$ nodes, with default connection parameters $\alpha=0.1$ and $\beta=\{0.03,0.02,0.015,0.01\}$ respectively, each with $500$ instances. We set these parameters so that the generated networks had similar node/link ratios as the real-world one (ASN). We denote the test sets as $\text{WA}^{\sf test}_{n}$ for node number $n$. The largest topology in training set had $4\times 10^{3}$ links, and the largest topology in test set had $2\times 10^{4}$ links, five times larger than training.
@@
B4 and ASN Datasets [^2] These are the smallest and largest topologies evaluated in [39] respectively, though we train and test on different SD pairs and demands than original paper.: Two real-world network topologies were used for evaluation which were also used in [39]: Google’s cloud-WAN topology B4 [18], and Microsoft’s AS-level Internet topology denoted as ASN [1]. B4 is a small topology with $12$ nodes and $38$ links. Thus we trained and tested on the same topology, with $200$ and $50$ training and test instances respectively. ASN is a large topology with $1739$ nodes and $8558$ links. For ASN, we chose two randomly sampled subgraphs as training set, with $217$ and $237$ nodes and each around $600$ links respectively; $2000$ instances were generated for each subgraph. We tested on the largest connected component of ASN after pruning all nodes in the training set, with $553$ nodes denoted as $\text{ASN}^{\sf test}_{553}$ and around $1600$ links, as well as the full ASN topology denoted as $\text{ASN}^{\sf test}_{1739}$; each had $500$ test instances. The largest test instance in ASN had more than 10×10$\times$10 × links than the largest training instance.
@@ V-B Implementation and Evaluation Setup
Comparison Algorithms: We compared TELGEN to two state-of-the-art learning-based TE solutions, TEAL [39] and HARP [3], as well as the optimal SciPy IPM solver as the baseline. Below, we describe their implementation details [^3] Our implementation code is available at https://github.ncsu.edu/NCSU-Yu-Research-Group/TE_TELGEN/.
@@ V-B1 TELGEN
We extended the GNN codebase in [7] to implement TELGEN. The GNN model consisted of MLP encoders each to encode the initial attributes of one set of vertices in $\{V^{\sf p},V^{\sf d}\cup V^{\sf l},\{v_{o}\}\}$, a series of graph convolutional (GCN) layers for message passing among vertices, and an MLP readout function to generate the TE solution from each path vertex’s attribute in $V^{\sf p}$; a ReLU layer was applied after readout to ensure non-negativity of the TE solution.
All encoders were two-layer MLPs with hidden dimensions $180$ and $360$. We used $J=2$ GCN layers for the inner loop. For ER and B4 datasets, we used $8$ outer layers; for Waxman and ASN, we used $16$ outer layers. The hidden dimensions of the readout two-layer MLP were $360$ and $720$. The model was trained independently for $150$ epochs on ER and WA datasets, and $50$ epochs on the B4 and ASN datasets.

### v2

@@ IV-B IPM Preliminaries
The primary bottleneck of IPM when solving large-scale LP problems is the complexity of solving large linear systems of equations (the Newton’s step in Line 5) at each iteration. The fastest linear system solver runs in almost cubic time, significantly limiting the scalability of IPM beyond modest problem sizes. Meanwhile, because of its use of the Newton’s step, IPM can converge to optimality in almost a constant number of iterations in most practical settings, typically less than 55-2020 iterations regardless of problem size [13]. This makes IPM a particularly good candidate to be mimicked and sped-up by ML—speeding-up the Newton’s step will yield good asymptotic improvement, while the small number of iterations is ideal for algorithmic alignment, as we discuss in Sec. IV-D.
@@ IV-C Representing TE as a Graph
To speed-up IPM, our first step is to properly represent the TE problem as a neural model that can be trained. As we motivated in Sec. IV-A, a simple GNN defined over the network topology graph itself cannot effectively embed information such as decision variables (path traffic split ratios), constraints and the objective. To this end, we develop a new GNN architecture, where the graph is not defined on the network but fully on the entities (variables, constraints and objective) and their correlations presented in the LP formulation in Program (2).
@@
Our approach differs from these works [43, 5, 33, 19] in two ways. First, [43] uses GNN to learn the final solution pattern directly. It does not encode the LP structure of TE, so its generalization is limited. Second, [5, 33, 19] apply GNNs to a selected subset of LPs not including TE. Formulating and solving TE requires specialized domain knowledge, for instance, edge-flow versus path-flow formulation, network-agnostic representation of variables and constraints, and reaching network generalizability goals. For [5], we tested and observed weak feasibility and poor generalizability to large network sizes due to simple GNN architecture design. We propose several key innovations, including path-flow formulation and neural representation of TE, dual-loop GNN architecture for generalizability, and TE-specific generalization-oriented training. We pretrain on small networks and evaluate on unseen instances. The model remains near-optimal on topologies up to $20\times$ larger (OGap $4\%$).
Figure 2: Double-looped GNN architecture: $K$ outer loops matching (expected) number of IPM iterations, and $J$ inner layers.
@@ IV-D Algorithmic Alignment and Strongly Supervised Training
Motivated by the algorithmic alignment in ML [5, 42], we want to better align the GNN model with the IPM algorithm to achieve strong performance and generalizability. Since an IPM iteration basically takes a Newton’s step from the previous iteration’s solution, starting from an initial feasible solution, we want the GNN model to learn the same behavior. To this end, we define a double-looped GNN architecture, which consists of $K$ outer GNN layers, and $J$ inner GNN layers in each outer layer, as in Fig. 2. Each outer layer corresponds to an IPM iteration, and seeks to mimic one Newton’s update on the previous layer’s solution (represented by vertex attributes after previous layer). The number of outer layers, $K$, should correspond to the number of iterations that IPM takes to converge, which is usually a constant that can be easily estimated [33].
Because each outer layer corresponds to the same operation (Newton’s step) that is applied on a different input (previous iteration’s solution), the message passing and aggregation functions $g^{k}$ and $MSG^{k}$ in all outer GNN layers should share the same neural network parameters. Furthermore, since the IPM generates a feasible solution in every iteration, each GNN outer layer’s output can be supervised by a loss function defined between the GNN’s output solution (by applying the readout function $r$) and the IPM’s updated solution. This strong supervision of each GNN layer’s output can significantly improve the model’s ability to learn not just the final optimal solutions in the training dataset, but also the trajectories of convergence to optimality by IPM. Such an ability is crucial for the neural model to learn the generalizability of the optimal algorithm—to converge to optimality given any problem instance.
One concern in the above architecture is whether one (outer) GNN layer can well approximate the Newton’s step, which is a fairly complex operation that warrants a nearly cubic complexity. To this end, we further expand each outer layer to consist of $J$ inner GNN layers. Intuitively, these layers are to learn the intermediate steps in solving a system of linear equations plus applying the system’s solution to the variables’ update. Since these steps do not correspond to feasible solutions to the original problem, the inner GNN layers are unsupervised. In practice, we find that $J=2$ inner layers are already sufficient for even the largest networks that we evaluate.
@@ IV-E Model Design and Training Procedures
With the above discussions, we now describe the GNN model and training procedures for solving large-scale TE problems with generalizability as depicted in Fig. 2. The GNN model consists of $J$ sets of message passing and aggregation functions $\bm{MSG}_{j}$ and $\bm{g}_{j}$ corresponding to the $J$ inner layers, and one readout function $r$. Each $\bm{MSG}_{j}$ contains six distinct functions: ${MSG}_{j}^{\sf p\text{-}dl}$, ${MSG}_{j}^{\sf dl\text{-}p}$, ${MSG}_{j}^{\sf p\text{-}o}$, ${MSG}_{j}^{\sf o\text{-}p}$, ${MSG}_{j}^{\sf dl\text{-}o}$, and ${MSG}_{j}^{\sf o\text{-}dl}$. Each $\bm{g}_{j}$ contains three functions: ${g}_{j}^{\sf p}$, ${g}_{j}^{\sf dl}$, and ${g}_{j}^{\sf o}$. For $\alpha,\beta\in\{{\sf p,dl,o}\}$, the functions ${MSG}_{j}^{\sf\alpha\text{-}\beta}$ and ${g}_{j}^{\sf\beta}$ are used to update the vertex attributes of $V^{\beta}$ based on the attributes of $V^{\alpha}$, where $\sf p$} denotes path vertices, $\sf dl$} denotes demand and link vertices, and $\sf o$} denotes the objective vertex. We cluster demand and link vertices together since all constraints can be normalized to have the same right-hand-side, i.e., let $\mathbf{b}=1$ in the LP formulation. All functions are shared across all outer layers, and the model can have a variable number of outer layers depending on the input instance.
@@
$h_{\text{dl}}^{k,j}=$ $g_{j}^{\text{dl}}\Big(h_{\text{dl}}^{k,j-1},MSG_{j}^{\text{p-dl}}(\{h_{\text{p}}^{k,j-1}|e_{p,i}\in E^{\text{pd}},e_{p,l}\in E^{\text{pl}}\}),$
$MSG_{j}^{\text{o-dl}}(\{h_{\text{o}}^{k,j-1}|e_{i,o}\in E^{\text{do}},e_{l,o}\in E^{\text{lo}}\})\Big);$
$h_{\text{o}}^{k,j}=$ $g_{j}^{\text{o}}\Big(h_{\text{o}}^{k,j-1},MSG_{j}^{\text{dl-o}}(\{h_{\text{dl}}^{k,j}|e_{i,o}\in E^{\text{do}},e_{l,o}\in E^{\text{lo}}\}),$
$MSG_{j}^{\text{p-o}}(\{h_{\text{p}}^{k,j-1}|e_{p,o}\in E^{\text{po}}\Big);$
$h_{\text{p}}^{k,j}=$ $g_{j}^{\text{p}}\Big(h_{\text{p}}^{k,j-1},MSG_{j}^{\text{dl-p}}(\{h_{\text{dl}}^{k,j}|e_{p,i}\in E^{\text{pd}},e_{p,l}\in E^{\text{pl}}\}),$
$MSG_{j}^{\text{o-p}}(\{h_{\text{o}}^{k,j}|e_{p,o}\in E^{\text{po}}\})\})\Big).$
Each outer iteration stacks $L$ graph-convolution layers aligned with IPM updates. To stabilize deep propagation and mitigate over-smoothing, we adopt residual connections [20, 41] in our model.
@@
where ReLU is a rectified linear unit, and $D^{(\tau)}$ and $L^{(\tau)}$ are the corresponding demand set and link set in problem instance $\tau$. While the IPM trajectory is feasible by construction, a learned GNN does not inherit these guarantees. Relying on the objective loss alone can produce infeasible solutions during intermediate iterations, which leads to either suboptimal convergence or infeasible final solutions. We therefore add an explicit feasibility signal at each outer loop. The constraint loss is defined as the constraint violation of solution generated in every GNN outer loop, normalized by the right-hand-side constant of each constraint. It penalizes the output solution of each outer loop to ensure feasibility of the solution update trajectory, in line with the original IPM.
@@
The objective loss penalizes when the GNN generates solutions that deviate too much from the IPM’s updates in each iteration.
@@
$\mathcal{L}=\rho_{1}\mathcal{L}_{\sf p}+\rho_{2}\mathcal{L}_{\sf dl}+\rho_{3}\mathcal{L}_{\sf o}.\vskip 3.61371pt$ (3)
@@ V-A Datasets
Erdös-Rényi (ER) Dataset: We generated random topologies based on the Erdös-Rényi model [9]. For training set, we generated networks with $\{20,30,\dots,100\}$ nodes, each with a connection probability of $\{0.3,0.4,\dots,0.8\}$. This resulted in $9\times 6=54$ topology configurations, and we generated $200$ instances for each configuration. For each instance, $10$ random SD pairs were selected. The link capacities and SD pair demands were both randomly generated in $[1000,5000]$; the same default range applies to all other datasets below. For test set, we generated significantly larger networks with $\{200,500,1000,2000\}$ nodes respectively, with connection probability in $\{0.1,0.2,\cdots,0.8,0.9\}$. Each of the 36 test configuration had $300$ instances with the same number of SD pairs and range for demands and capacities; we denote these $36$ test sets as $\text{ER}^{\sf test}_{n,q}$, where $n$ is the node number, $q$ is the connection probability. The largest topology in training set had almost $8\times 10^{3}$ links, and the largest topology in test set had $3.6\times 10^{6}$ links, two to three orders of magnitude larger than training.
@@
Waxman (WA) Dataset: Random topologies were generated based on the Waxman model [38]. For training set, we generated networks with $\{200,300,\dots,800\}$ nodes with the default connection parameters $\alpha=0.1$ and $\beta=\{0.2,0.18,\dots,0.08\}$ [^1] Here the $\alpha,\beta$ are Waxman parameters [38], not those used in Sec. IV-E.. Each configuration had $800$ instances, and each instance had $20$ random SD pairs. For test set, we generated large networks that had $\{2000,3000,4000,5000\}$ nodes, with the default connection parameters $\alpha=0.1$ and $\beta=\{0.03,0.02,0.015,0.01\}$ respectively, each with $500$ instances. We set these parameters so that the generated networks have similar node/link ratios as the real-world one (ASN). We denote the test sets as $\text{WA}^{\sf test}_{n}$ for node number $n$. The largest topology in training set had $4\times 10^{3}$ links, and the largest topology in test set had $2\times 10^{4}$ links, five times larger than training.
@@
B4 and ASN Datasets: Two real-world network topologies were used for evaluation which were also used in existing work [43]: Google’s cloud-WAN topology B4 [16], and Microsoft’s AS-level Internet topology denoted as ASN [35]. B4 is a small topology with $12$ nodes and $38$ links. Thus we trained and tested on the same topology, with $200$ and $50$ training and test instances respectively. ASN is a large topology with $1739$ nodes and $8558$ links. For ASN, we chose two randomly sampled subgraphs as training set, with $217$ and $237$ nodes and each around $600$ links respectively; $2000$ instances were generated for each subgraph. We tested on the largest connected component of ASN after pruning all nodes in the training set, with $553$ nodes denoted as $\text{ASN}^{\sf test}_{553}$ and around $1600$ links, as well as the full ASN topology denoted as $\text{ASN}^{\sf test}_{1739}$; each had $500$ test instances. The largest test instance in ASN had more than $10\times$ links than the largest training instance.
@@ V-B Implementation and Evaluation Setup
Comparison Algorithms: We compared TELGEN to two state-of-the-art learning-based TE solutions, TEAL [43] and HARP [2], and also the optimal SciPy IPM solver as the baseline. [^2] In addition to these, we have also tested Google’s OR-Tools [29] open-source LP solver. Unfortunately, OR-Tools cannot generate feasible solutions consistently as the other solvers, and fails to solve many instances when network size becomes large (e.g. $n\geq 1000$). Hence we omit the comparison results with OR-Tools below. Below, we describe their implementation details.
@@ V-B1 TELGEN
We extended the GNN codebase in [6] to implement our model. The GNN model was composed of MLP encoders each to encode the initial attributes of one set of vertices in $\{V^{\sf p},V^{\sf d}\cup V^{\sf l},\{v_{o}\}\}$, a series of graph convolutional (GCN) layers to perform messaging among vertices, and an MLP readout function to generate the TE solution from each path vertex’s attribute in $V^{\sf p}$; a ReLU layer was applied after the readout function to ensure non-negativity of the TE solution.
All encoders were all two-layer MLPs with hidden dimensions 180 and 360. We used $J=2$ GCN layers for the inner loop. For ER and B4 datasets, we used $8$ outer layers; for Waxman and ASN, we used $16$ outer layers. The hidden dimensions of the readout two-layer MLP were 360 and 720. The model was trained independently for $150$ epochs on ER and WA datasets, and $50$ epochs on the B4 and ASN datasets.